OpenStudy (anonymous):

Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form

OpenStudy (saifoo.khan):

perpendicular slope = \[4 \times m_2 = -1\] \[m_2 = -\frac14\]

OpenStudy (saifoo.khan):

slope = -1/4 point= (2,-3)

OpenStudy (anonymous):